OpenStudy (anonymous):

Non Homogeneous Linear equations \[y''+4y=e^{3x}\] I understand that the auxiliary equation is \[r^2+4=0\] The roots are \[\pm 2\] but this is about as for as I understand the process

OpenStudy (anonymous):

Why do we have sines and cosines or is that just part of the process?

OpenStudy (anonymous):

+/- 2i

OpenStudy (anonymous):

which means use a linear combination of sin and cos for the for of your solution

OpenStudy (anonymous):

constants go out front: C1*sin ( ) eigenvalues are the angular freq: C1*sin(2*t)

OpenStudy (anonymous):

whoops I left a bit out solutions should actually look like : C1*cos(2*t) + C2*i*sin(2*t) (I left out the 'i')

OpenStudy (anonymous):

that help?

OpenStudy (anonymous):

oh I remember now

OpenStudy (anonymous):

to get a particular solution of the non homogeneous equation \[ y= A e ^{3 x}\\ y'= 3 A e^{3 x}\\ y''=9 A e^{3 x}\\ y''(x) + 4 y[x]= 13 Ae^{3 x}=e^{3 x}\\ A=\frac 1 {13} y(x)=\frac 1 {13}e^{3 x} \] The general solution is \[ y= C_1 \cos(2x) + c_2 \sin(2x) +\frac 1 {13}e^{3 x} \]